One side of a square was decreased by 3. The other side stayed the same . The area of the neve rectangle is 75% of the area of the square. Write and solve an equation to find the side of the original square.
Answer by ikleyn(52781)

.

x*(x-3) = 0.75x^2


Since x is not zero, you can divide both sides by x. You will get


x - 3 = 0.75x

x - 0.75x = 3

0.25x = 3

x = 3/0.25 = 12.


ANSWER.  The side length of the original square is 12 units.

Solved.